# Sharp Self-Normalized Concentration Inequalities of Marginal Mean with Sample Variance Only

### v2

@@ 2.2.2 Inequalities
For each given $n\geq 1$, let $\mathcal{H}_{n}:=\{h_{n}:\mathcal{Z}\to\mathbb{R}\}$ is a class of measurable mappings such that $h_{n}\in[a_{n},b_{n}]\subset\mathbb{R}$ holds for any given $n$ and $\mathbb{E}[h_{n}(Z_{i})]=\mu_{n}$ holds for any given $i$ and $n$. As for the dependence, we majorly consider the cases where process $\mathbf{Z}$ satisfies $\phi$ or $\tilde{\phi}-$mixing conditions introduced as follow. By letting $\mathbb{P}_{n}(h_{n})=\frac{1}{n}\sum_{i=1}^{n}h_{n}(Z_{i})$, this main goal of section is to give self normalized concentration inequality of $|\mathbb{P}_{n}(h_{n})-\mu|$ for $\phi/\tilde{\phi}-$mixing processes without assuming summability.
@@
$\mathbb{P}\left(\Big|\frac{1}{n}S_{nn}(h_{n})\Big|\leq\tilde{\nu}_{n}(\delta)\max\{A_{\phi},B_{\phi}\}+\frac{[l]}{n}(b_{n}-a_{n})\right)\geq 1-2\delta$
@@
$\tilde{\nu}_{n}(\delta)=\Big(1-\sqrt{\frac{2\log(\frac{1}{\delta})}{m}}\Big)^{-1},B_{\phi}=\frac{2([l]+2\Phi_{n})(b_{n}-a_{n})}{n},$
@@
$\mathbb{P}\left(\Big|\frac{1}{n}S_{nn}(h_{n})\Big|\leq\tilde{\nu}_{n}(\delta)\max\{A_{\tilde{\phi}},B_{\tilde{\phi}}\}+\frac{[l]}{n}(b_{n}-a_{n})\right)\geq 1-2\delta,$
@@
Theorems 2 and 3 extend the martingale-difference bound of Theorem 1 to weakly dependent time series. Under mild $\phi/\tilde{\phi}-$mixing assumptions, we obtain self-normalized Bernstein-type inequalities in which the variance term is a block-based empirical long-run variance $\hat{V}_{n}$ that is fully observable from the data. The structure of the bounds is exactly analogous to the martingale case, in which the leading self-normalized term $\sqrt{\frac{2\hat{V}_{n}}{n}\log(\frac{1}{\delta})}$ and a coverage probability $1-\delta$. In particular, we do not impose or use any explicit decay rate for the mixing coefficients. Mixing coefficient affects the high-probability bound in a very weak way. Conceptually, these results provide fixed-$n$ empirical Bernstein inequalities for the marginal mean of weakly dependent sequences, with a variance scale that automatically adapts to the appropriate long-run variance whenever it exists. This contrasts with classical Bernstein-type inequalities for mixing processes, which typically involve non-empirical long-run variance bounds and explicit mixing coefficients, or pay a loss in rate due to blocking. Here the only variance term is the simple block-based empirical long-run variance, and the resulting bounds retain a parametric $\sqrt{n}$-rate and the same sharp constant $\sqrt{2\log(\frac{1}{\delta})}$ as in the martingale-difference case. The knowledge of mixing coefficient affects the construction high-probability bound in a very indirect way. For example, under many circumstances, it is natural to assume the mixing coefficients are summable (i.e. $\sum_{k=1}^{\infty}\tilde{\phi}(k)<\infty$, $\sum_{k=1}^{\infty}{\phi}(k)<\infty$). Thus, replacing them with any user-defined diverging sequence, like $\log\log n$, yield a valid high-probability bound and the sharpness is not harmed at all since leading term is still $\sqrt{\frac{2\hat{V}_{n}}{n}\log(\frac{1}{\delta})}$.
Unfortunately, we sometimes have to face some “strongly-dependent” situations where the assumption of summability of mixing coefficients may NOT hold, like the case where we can only assume $\tilde{\phi}(k)\lesssim 1/k$. Under this circumstance, we can not even guarantee the summability of autocovariances and the existence of the classical long-run variance. A more severe question is that it is often difficult to detect diverging speed of $\sum_{k=1}^{n}\tilde{\phi}(k)=:\tilde{\Phi}_{n}$ and $\sum_{k=1}^{n}\phi(k)=:\Phi_{n}$. Hence, according to our discussion in Section 2.1, we can get out of this awkward situation by ignoring the terms containing unknown parameters. The following Theorem 4 precisely describe the cost of this ignorance. Similar to Corollary 2, the cost is also on an exponential-level, which is usually affordable.
@@
Suppose $\mathbf{Z}$ is a $\mathcal{Z}$-valued $\phi/\tilde{\phi}-$mixing process. Based on the notations used in Theorems 2 and 3, for all $\delta>0$, $n\geq 1$ and user-defined $\theta^{\prime}\in(0,\frac{1}{2})$, $c_{n}\searrow 0$ and $t_{n}\searrow 0$, we have
$\mathbb{P}\Big(|\frac{1}{n}S_{nn}(h_{n})|\leq\tilde{\nu}_{n}(\delta)\max\{\mathbf{A}^{\prime},\mathbf{B}^{\prime}\}+c_{n}\Big)\geq 1-2\delta-\sum_{k=1}^{3}\text{Error}_{k},$
where $\mathbf{A}^{\prime}=\sqrt{2\frac{1}{n}\log(\frac{1}{\delta})\hat{V}_{n}}+t_{n}\sqrt{2\log(\frac{1}{\delta})}+\sqrt{\frac{2\xi_{n}\log(\frac{1}{\delta})}{n^{2}}},$
$\mathbf{B}^{\prime}=\frac{2\log(\frac{1}{\delta})[l](b_{n}-a_{n})}{n}+\frac{\log(\frac{1}{\delta})}{n^{1-\theta^{\prime}}},\ \text{Error}_{1}=2\exp\left(-2\Big(\frac{(n/[l])c_{n}}{(b_{n}-a_{n})}\Big)^{2}\right),$
$\text{Error}_{2}=2m\exp\left(-0.5\Big(\frac{\sqrt{n[l]}t_{n}}{\triangle_{n}\Psi_{n}}\Big)^{2}\right),\ \text{Error}_{3}=2m\exp\left(-\Big(\frac{n^{\theta^{\prime}}}{16\triangle_{n}\Psi_{n}}\Big)^{2}\right).$
Here $\Psi_{n}$ indicates $\Phi_{n}$ and $\widetilde{\Phi}_{n}$ under $\phi$ and $\tilde{\phi}-$mixing conditions respectively. $\triangle_{n}$ means $2(b_{n}-a_{n})$ and $h_{n}||_{TV}$ under $\phi$ and $\tilde{\phi}-$mixing conditions respectively
Suppose $a_{n}=a$ and $b_{n}=b$. For $\phi$-mixing process with coefficient $\phi(k)=\frac{1}{k}$, the ignorance of parameters $\Phi_{n}$ is nearly harmless. Generally speaking, even in regimes where the autocovariances are not summable and a classical long-run variance may fail to exist, Theorem 4 still delivers stable, fully data-driven high-probability interval based on the block empirical variance. In such cases we may not have any CLT-type interpretation, but the bounds remain valid and operational, providing finite-sample control of the marginal mean under very strong dependence.
@@ 3 Conclusions
In this paper we have developed a family of self-normalized concentration inequalities for marginal means, in both martingale-difference and weakly dependent settings, with variance terms that are fully observable from the data. In the martingale-difference (and in particular i.i.d.) case, our bounds take the form of empirical Bernstein inequalities where the self-normalizer is the usual sample variance. In the $\phi/\tilde{\phi}-$mixing case, the self-normalizer becomes a simple block-based empirical long-run variance. In all cases, the bounds have an explicit Freedman/Bernstein-type structure: a leading term of order $\sqrt{\text{empirical (long run) variance}\log(\frac{1}{\delta})/n}$ and a soft floor of order $\frac{b}{n}\log(\frac{1}{\delta})$, with coverage probability $1-\delta$ up to an exponentially small remainder.
We view these results as a foundational step towards a broader theory of sharp, fully empirical self-normalized inequalities for dependent data. A more detailed follow-up, including uniform-in-time versions, statistical applications, and high-dimensional extensions, is in preparation.
@@ A.1 Technical Lemmas
Given any two positive integers $m$ and $l$, define $\{1,...,ml\}=\bigcup_{j=1}^{m}B_{j}$, where $B_{j}=\{(j-1)l+1,...,jl\}$. Suppose $\{Z_{i}\}_{i=1}^{ml}$ is a group of real numbers. By denoting $\bar{w}=\frac{1}{ml}\sum_{i=1}^{ml}Z_{i}$, we have the following identity holds for every $m,l\geq 1$,
$\sum_{j=1}^{m}(\sum_{i\in B_{j}}(Z_{i}-\bar{w}))^{2}+l^{2}m(\bar{w}-\mu)^{2}=\sum_{j=1}^{m}(\sum_{i\in B_{j}}(Z_{i}-\mu))^{2}$
@@
$\sigma(\mathcal{R})=\sigma(\mathcal{A}\cup\sigma(\Lambda))=:\mathcal{A}\lor\sigma(\Lambda),$
where $\sigma(\mathcal{R})$ and $\sigma(\Lambda)$ are the sigma algebras generated by set $\mathcal{R}$ and random variable $\Lambda$.
@@
Suppose $(\Omega,\mathcal{F},\mathbb{P})$ is a probability space and $\mathcal{A}\subset\mathcal{F}$ is a sub $\sigma$-algebra. Suppose $\Lambda$ is a random variable taking value in Borel measurable space $(E,\mathcal{B}(E))$. Provided that $\phi:E\times\Omega\to[0,\infty)$ is measurable with respect to $\mathcal{B}(E)\otimes\mathcal{A}$,
@@
$\tilde{g}_{K}(\lambda,\omega)$ is obviously measurable with respect to $\mathcal{B}(E)\times\mathcal{A}$. Then, for any non-negative mapping $h$ measurable with respect to $\mathcal{B}(E)\otimes\mathcal{A}$, there exists a sequence of non-negative simple functions $\{h_{K}\}$ such that (i) $h_{K}\leq h_{K+1}$; (ii) $\lim_{K\to\infty}h_{K}=h$ pointwisely. Then, for any given $\lambda\in E$ and $\omega\in\Omega$,

### v3

@@ 2.2.2 Inequalities
For each given $n\geq 1$, let $\mathcal{H}_{n}:=\{h_{n}:\mathcal{Z}\to\mathbb{R}\}$ is a class of measurable mappings such that $h_{n}\in[a_{n},b_{n}]\subset\mathbb{R}$ holds for any given $n$ and $\mathbb{E}[h_{n}(Z_{i})]=\mu_{n}$ holds for any given $i$ and $n$. As for the dependence, we primarily consider the cases where process $\mathbf{Z}$ satisfies $\phi$ or $\tilde{\phi}-$mixing conditions introduced as follow. By letting $\mathbb{P}_{n}(h_{n})=\frac{1}{n}\sum_{i=1}^{n}h_{n}(Z_{i})$, this main goal of section is to give self normalized concentration inequality of $|\mathbb{P}_{n}(h_{n})-\mu|$ for $\phi/\tilde{\phi}-$mixing processes without assuming summability.
@@
$\mathbb{P}\left(\Big|\mathbb{P}_{n}(h_{n})-\mu\Big|\leq\tilde{\nu}_{n}(\delta)(A_{\phi}+B_{\phi})+\frac{n-m[l]}{n}(b_{n}-a_{n})\right)\geq 1-3\delta$
@@
$\tilde{\nu}_{n}(\delta)=\Big(1-\sqrt{\frac{2\log(\frac{1}{\delta})}{m}}\Big)^{-1},B_{\phi}=\frac{(3.15[l]+2\Phi_{n})(b_{n}-a_{n})\log(\frac{1}{\delta})}{n},$
@@
$\mathbb{P}\left(\Big|\mathbb{P}_{n}(h_{n})-\mu\Big|\leq\tilde{\nu}_{n}(\delta)(A_{\tilde{\phi}}+B_{\tilde{\phi}})+\frac{n-m[l]}{n}(b_{n}-a_{n})\right)\geq 1-3\delta,$
@@
Theorems 2 and 3 extend the martingale-difference bound of Theorem 1 to weakly dependent time series. Under mild $\phi/\tilde{\phi}-$mixing assumptions, we obtain self-normalized Bernstein-type inequalities in which the variance term is a block-based empirical long-run variance $\hat{V}_{n}$ that is fully observable from the data. The structure of the bounds is exactly analogous to the martingale case, in which the leading self-normalized term $\sqrt{\frac{2\hat{V}_{n}}{n}\log(\frac{1}{\delta})}$, a linear term with slightly enlarged constant and a coverage probability $1-3\delta$. In particular, we do not impose or use any explicit decay rate for the mixing coefficients. Mixing coefficient affects the high-probability bound in a very weak way. Conceptually, these results provide fixed-$n$ empirical Bernstein inequalities for the marginal mean of weakly dependent sequences, with a variance scale that automatically adapts to the appropriate long-run variance whenever it exists. This contrasts with classical Bernstein-type inequalities for mixing processes, which typically involve non-empirical long-run variance bounds and explicit mixing coefficients, or pay a loss in rate due to blocking. Here the only variance term is the simple block-based empirical long-run variance, and the resulting bounds retain a parametric $\sqrt{n}$-rate. The knowledge of mixing coefficient affects the construction high-probability bound in a very indirect way. For example, under many circumstances, it is natural to assume the mixing coefficients are summable, which means $\max_{n}\Phi_{n}\lor\tilde{\Phi}_{n}<\infty$. Thus, replacing them with any user-defined diverging sequence, like $\log\log n$, yield a valid high-probability bound and the sharpness is not harmed at all since leading term is still $\sqrt{\frac{2\hat{V}_{n}}{n}\log(\frac{1}{\delta})}$.
Unfortunately, there are some “strongly-dependent” situations where the assumption of summability (e.g. $\sum_{k=1}^{\infty}\tilde{\phi}(k)<\infty$) of mixing coefficients may NOT hold, like the case where we can only assume $\tilde{\phi}(k)\lesssim 1/k$. Under this circumstance, we can not even guarantee the summability of auto-covariances and the existence of the classical long-run variance. A more severe question is that it is often difficult to detect diverging speed of $\sum_{k=1}^{n}\tilde{\phi}(k)=:\tilde{\Phi}_{n}$. According to our discussion in Section 2.1, it seems that we can get out of this awkward situation by ignoring the terms containing this unknown parameter. The following Theorem 4 precisely describe the cost of this ignorance. Similar to Corollary 2, the cost is also on an exponential-level, which is usually affordable.
@@
Suppose $\mathbf{Z}$ is a $\mathcal{Z}\subset\mathbb{R}$-valued $\tilde{\phi}-$mixing process. Based on the notations used in Theorems 2 and 3, for all $\delta>0$, $n\geq 1$ and user-defined $c_{n},t_{n},s_{n}>0$, when $n\neq m[l]$,
$\ \ \ \ \ \ \ \ \ \ \ \mathbb{P}(|\mathbb{P}_{n}(h_{n})-\mu|\leq\tilde{\nu}_{n}(\delta)U_{n}^{\prime}(\delta))\geq 1-3\delta-\sum_{k=1}^{3}\text{Error}_{k},$
$U_{n}^{\prime}(\delta)$ $=\sqrt{\frac{2\log(\frac{1}{\delta})\hat{V}_{n}}{n}}++(1+\frac{1}{n})t_{n}\sqrt{2\log(\frac{1}{\delta})}+3.15\log(\frac{1}{\delta})\Big(\frac{l(b_{n}-a_{n})}{n}+s_{n}\Big)+c_{n},$
[…]
When $m[l]=n$, the result above holds by letting $c_{n}=\text{Error}_{1}=0$.
Suppose $h_{n}||_{TV}=O(1)$. For $\tilde{\phi}-$mixing process with coefficient $\tilde{\phi}(k)=\frac{1}{k}$, the ignorance of parameters $\tilde{\Phi}_{n}:=\sum_{k=1}^{n}\tilde{\phi}(k)$ is nearly harmless. Generally speaking, even in regimes where the auto-covariances are not summable and a classical long-run variance may fail to exist, Theorem 4 still delivers stable, fully data-driven high-probability interval based on the block empirical variance. In such cases we may not have any CLT-type interpretation, but the bounds remain valid and operational, providing finite-sample control of the marginal mean under very strong dependence.
@@ 3 Conclusions
This paper develops a simple route to empirical Bernstein–type concentration for dependent data. For bounded martingale difference sequences, we obtain a one–parameter Freedman–style inequality with predictable quadratic variation, and then show how to pass from the predictable quadratic variation to the empirical quadratic variation. This yields a fully empirical self–normalized bound in which the variance term is entirely data–driven. For $\phi/\tilde{\phi}-$mixing sequences, we combine a blocking scheme with the martingale results to derive analogous empirical Bernstein inequalities based on block empirical long–run variance, without imposing explicit decay rates on the mixing coefficients.
A technical ingredient underlying our approach is a conditional measurability mechanism, formalized in Lemmas 3 and 4, which allows us to rigorously justify the use of random, data–dependent tilting parameters in exponential supermartingale arguments. This mechanism is independent of any particular model and may be of separate interest for other self–normalized or “empirical variance” inequalities under dependence.
The present paper focuses on the core probabilistic tools and their basic forms. In ongoing work, we apply these inequalities to concrete statistical procedures—such as adaptive inference and sequential methods under temporal dependence—showing that the empirical normalization leads to practical, fully data–driven uncertainty quantification without prior knowledge of long–run variance.
4 Acknowledgement
The author is grateful to Prof. Dr. Aaditya Ramdas for pointing out an important mistake in the earlier version of this work.
@@ A.1 Technical Lemmas
Given any two positive integers $m$ and $l$, define $\{1,...,ml\}=\bigcup_{j=1}^{m}B_{j}$, where $B_{j}=\{(j-1)l+1,...,jl\}$. Suppose $\{Z_{i}\}_{i=1}^{ml}$ is a group of real numbers. By denoting $\bar{w}=\frac{1}{ml}\sum_{i=1}^{ml}Z_{i}$, we have the following identity holds for all $\mu\in\mathbb{R}$ and $m,l\geq 1$,
$\sum_{j=1}^{m}(\sum_{i\in B_{j}}(Z_{i}-\bar{w}))^{2}+ml^{2}(\bar{w}-\mu)^{2}=\sum_{j=1}^{m}(\sum_{i\in B_{j}}(Z_{i}-\mu))^{2}\ \ \ \ \text{a. s.}$
@@
$\sigma(\mathcal{R})=\sigma(\mathcal{A}\cup\mathcal{B})=:\mathcal{A}\lor\mathcal{B}.$
@@
Suppose $(\Omega,\mathcal{F},\mathbb{P})$ is a probability space and $\mathcal{A}\subset\mathcal{F}$ is a sub $\sigma$-algebra. Suppose $\Lambda$ is a random variable taking value in standard Borel measurable space $(E,\mathcal{B}(E))$. Provided that $\phi:E\times\Omega\to[0,\infty)$ is measurable with respect to $\mathcal{B}(E)\otimes\mathcal{A}$,
@@
Lemma A.7
((6.2) in ? (?)) Suppose $\mathbf{Z}:=(Z_{i})_{i\in\mathbb{Z}}$ is a $\tilde{\phi}-$mixing process. Based on the $h_{n}$ introduced in Theorem 3, the following inequality holds for every given $m\geq 1$ and $t>0$,
[…]
@@
$\tilde{g}_{K}(\lambda,\omega)$ is obviously measurable with respect to $\mathcal{B}(E)\times\mathcal{A}$. Then, for any non-negative mapping $h$ measurable with respect to $\mathcal{B}(E)\otimes\mathcal{F}$, there exists a sequence of non-negative simple functions $\{h_{K}\}$ such that (i) $h_{K}\leq h_{K+1}$; (ii) $\lim_{K\to\infty}h_{K}=h$ pointwisely. Then, for any given $\lambda\in E$ and $\omega\in\Omega$,